OpenStudy (anonymous):

sin(feta)=-.06514 (0 degress less than or equal to feta <360)

OpenStudy (anonymous):

\[180+\frac{180 \sin ^{-1}(0.06514)}{\pi }=183.735 \text{ degrees} \]

OpenStudy (campbell_st):

isn't there also a 4th quadrant solution to this question... sine is negative in the third quadrant where angles are between 180 and 270 degrees. and negative in the 4th quadrant where angles are between 270 and 360 degrees so find the value of \[\theta = \sin^{-1}(0.06514)\] then the 3rd quadrant solution is \[180 + \theta\] 4th quadrant solution \[360 - \theta\] hope this helps.

OpenStudy (anonymous):

thanks it did :)